OpenStudy (anonymous):

\[\int\limits_{}((logx)/(x^{3}))dx\]

OpenStudy (jamesj):

So integrate by parts with u = ln x and v' = 1/x^3

OpenStudy (turingtest):

is it a natural log?

OpenStudy (anonymous):

yahh

OpenStudy (turingtest):

then do what james said

OpenStudy (anonymous):

I would prefer JamesJ's approach

OpenStudy (jamesj):

what have you written? That's not right.

OpenStudy (jamesj):

\[ \int uv' = uv - \int u'v \] with \( u = \ln x, \ v = 1/x^3 \).

OpenStudy (jamesj):

@amogh, again, no.

OpenStudy (jamesj):

@korcan, I suggest you just try and solve the problem yourself and let us know if you get stuck.

OpenStudy (anonymous):

x.O @ amogh

OpenStudy (anonymous):

Oh...I'm very sorry! Lost myself. I'll check

OpenStudy (anonymous):

OK, I am trying it ,

OpenStudy (anonymous):

@korkankonaglu: Don't mind my answer, that's not right

OpenStudy (jamesj):

correction to my last math post, u = ln x, v' = 1/x^3. Not v = 1/x^3

OpenStudy (anonymous):

ohh k , what was v ?

OpenStudy (jamesj):

that's for you to find.

OpenStudy (jamesj):

\[ \int uv' = uv - \int u'v \] with \[ u = \ln x, \ v' = 1/x^3 \]

OpenStudy (anonymous):

v' = dv/dx , if that's interrupting you

OpenStudy (anonymous):

here, I think I can use it: \[\int\limits_{}p(x)logxdx = Q(X)logx - \int\limits_{}\S(x)dx\] I mean if we take p(x) as x^-3, then we will have : x^-3 * logx * dx and by using it: we will have : (-1/(2x^2))logx - (1/(4x^2)) + C ,,,,,,,,,,,,, \[\int\limits_{}P(x)logxdx ,, u=logx, P(x)dx=dv\] \[du=(dx)/x, v=\int\limits_{}P(x)dx = Q(x)\] q(x), is a polnomial of (n+1)degrees. by these all then we will have : \[\int\limits_{}P(x)logxdx = Q(x)logx - \int\limits_{}\S(x)dx\] if I have to give an example: \[\int\limits_{}x^n logxdx = ? \] u = logx, x^ndx = dv du = dx/x , x^((n+1))/ (n+1) = v then : \[\int\limits_{}x^n logxdx = (x^{n+1} / (n+1)) - \int\limits_{}(x^n / (n+1))dx\] and etc.... IS IT RIGHT? :D

OpenStudy (anonymous):

Hello guys is it right?

OpenStudy (jamesj):

yes, looks right.